OpenStudy (anonymous):

Translate the word phrase into a equation. seven times the difference of a number and 3 is 35. I am stuck on this!

OpenStudy (dangerousjesse):

Difference of = the numbers are already in their respective orders

OpenStudy (dangerousjesse):

So \(7\times x-3=35\)

OpenStudy (dangerousjesse):

Or just 7x :)

OpenStudy (anonymous):

i would think maybe \[7(x-3)=35\]

OpenStudy (anonymous):

thats choice C lol

OpenStudy (dangerousjesse):

Satellite is correct, I always manage to miss something :P

OpenStudy (anonymous):

that okay thank u guys! :)